# ProxyLM: Predicting Language Model Performance on Multilingual Tasks via Proxy Models

### v2

@@ Appendix C Feature Importances
We provide feature importance scores of XGBoost with ProxyLM Ensemble for the random English-centric experiment in Figure 15 and 16. Each combination consists of one most influential feature followed by others with marginal contributions to the model, each with an importance score of 0.12 or less. We observe that proxy models are always the most influential features in prediction.
Appendix D License for Artifacts
@@
Hyper-parameter Many-to-Many Langs. English-centric
@@
min_child_weight 5 2.5 3 3.5
max_depth 3 3 5 4
gamma 0 0 0.2 0.1
subsample 0.7 0.9 0.95 0.8
colsample_bytree 0.6 0.6 0.85 1.0
reg_alpha 0 0 0.1 0.1
reg_lambda 0.35 0.15 0.1 0.3
Table 14: List of hyper-parameters used for XGBoost Regressor with M2M100 and NLLB models trained with English-centric and Many-to-Many Languages datasets.
@@
Table 15: List of hyper-parameters used for Poly2/Poly3 Regressor with M2M100 and NLLB models trained with English-centric and Many-to-Many Languages datasets.
@@
Table 16: List of hyper-parameters used for LGBM Regressor with M2M100 and NLLB models trained with English-centric and Many-to-Many Languages datasets. “Max" indicates the maximum value set for the hyper-parameter during the hyper-parameter search.
@@
Table 17: List of hyper-parameters used for MF Regressor with M2M100 and NLLB models trained with English-centric and Many-to-Many Languages datasets. “Max" indicates the maximum value set for the hyper-parameter during the hyper-parameter search.
@@
Random (English-centric) 829 356
@@
Unseen 1,084 101
Cross-Dataset 1,185 224
Table 18: Regressor’s training and test set size on different experimental settings. The total experimental records for English-centric and Many-to-Many Languages datasets are 1,185 and 224, respectively.
@@
FLoRes Costa-jussà et al. (2022b) Muennighoff/flores200 CC-BY-SA 4.0
@@
Figure 5: Detailed results of XGBoost with ProxyLM Ensemble on NLLB model under the LOLO setting using the English-centric dataset from Table 1. The results are grouped by (a) Joshi Class and (b) language family that follows the mapping which is provided in Appendix A.1; (c) shows the scatter plot illustrating the correlation of spBLEU scores between the ProxyLM’s prediction and estimated LM, with the gray dashed line representing the line of equality ($y=x$) with $R^{2}=0.89$ and black dashed line representing Locally Weighted Scatterplot Smoothing (LOWESS) curve to represent the trend.
Figure 6: Detailed results of XGBoost with ProxyLM Ensemble under the LOLO setting using the English-centric dataset from Table 1, grouped by vitality, for (a) M2M100 and (b) NLLB model. Vitality is outlined in Table 8.
Figure 7: Detailed results of XGBoost with ProxyLM Ensemble on the M2M100 model under the LOLO setting using the English-centric dataset from Table 1, categorized by language.
Figure 8: Performance heatmap of XGBoost with ProxyLM Ensemble on the M2M100 model under the LOLO setting using the English-centric dataset from Table 1. The heatmap showcases the performance based on vitality level and language family, as outlined in Table 8.
Figure 9: Detailed results of XGBoost with ProxyLM Ensemble on the NLLB model under the LOLO setting using the English-centric dataset from Table 1, categorized by language.
Figure 10: Performance heatmap of XGBoost with ProxyLM Ensemble on the NLLB model under the LOLO setting using the English-centric dataset from Table 1. The heatmap illustrates the model’s performance based on vitality levels and language family classifications as specified in Table 8.
Figure 11: Detailed results of XGBoost with ProxyLM Ensemble on the M2M100 model under the LOLO setting using the Many-to-Many Languages dataset from Table 3, organized by language groupings. The mapping of (a) Vitality, (b) Joshi Class, and (c) language family align with the classifications provided in Table 9.
Figure 12: Detailed results of XGBoost with ProxyLM Ensemble on the M2M100 model under the LOLO setting using the Many-to-Many Languages dataset from Table 3, categorized by language.
Figure 13: Detailed results of XGBoost with ProxyLM Ensemble on the NLLB model under the LOLO setting using the Many-to-Many Languages dataset from Table 3, organized by language groups. The mapping of (a) Vitality, (b) Joshi Class, and (c) language family corresponds to the classifications provided in Table 9.
Figure 14: Detailed results of XGBoost with ProxyLM Ensemble on the NLLB model under the LOLO setting using the Many-to-Many Languages dataset from Table 3, categorized by language.
Figure 15: Feature importance analysis of XGBoost with ProxyLM Ensemble on the M2M100 model using the English-centric dataset.
Figure 16: Feature importance analysis of XGBoost with ProxyLM Ensemble on the NLLB model using the English-centric dataset.

### v3

@@ Appendix D Feature Importances
We provide feature importance scores of XGBoost with ProxyLM Ensemble for the random English-centric experiment in Figure 8 and 9. Each combination consists of one most influential feature followed by others with marginal contributions to the model, each with an importance score of 0.12 or less. We observe that proxy models are always the most influential features in prediction. Other feature importances plot are also provided in Figure 11 until 13.
Appendix E License for Artifacts
@@
Hyper-parameter Using spBLEU Using COMET
@@
min_child_weight 5.0 4.2 3.2 1.1
max_depth 5 4 3 5
gamma 0 0 0 0
subsample 0.6 0.94 0.6 1
colsample_bytree 0.83 0.82 0.9 0.86
reg_alpha 0.2 0.32 0.11 0
reg_lambda 0.1 0.37 0.48 0.05
Table 21: List of hyper-parameters used for XGBoost Regressor on MT task with M2M100 and NLLB models trained with English-centric dataset.
Hyper-parameter M2M100 NLLB
max n_estimators 2000 2000
max eta 0.1 0.1
min_child_weight 5 2.5
max_depth 3 3
gamma 0 0
subsample 0.7 0.9
colsample_bytree 0.6 0.6
reg_alpha 0 0
reg_lambda 0.35 0.15
Table 22: List of hyper-parameters used for XGBoost Regressor on MT task with M2M100 and NLLB models trained with Many-to-Many Languages dataset.
Hyper-parameter Aya-23 LLaMA3
max n_estimators 5000 5000
max eta 0.1 0.1
min_child_weight 3.0 3.0
max_depth 3 3
gamma 0.1 0.1
subsample 0.85 0.6
colsample_bytree 1.0 0.95
reg_alpha 0.1 0.1
reg_lambda 0.2 0.5
Table 23: List of hyper-parameters used for XGBoost Regressor on intent classification and slot filling tasks with Aya-23 and LLaMA-3 models.
@@
Table 24: List of hyper-parameters used for Poly2/Poly3 Regressor for all tasks.
@@
Table 25: List of hyper-parameters used for LGBM Regressor for all tasks. “Max" indicates the maximum value set for the hyper-parameter during the hyper-parameter search.
@@
Table 26: List of hyper-parameters used for MF Regressor with M2M100 and NLLB models trained with Many-to-Many Languages datasets. “Max" indicates the maximum value set for the hyper-parameter during the hyper-parameter search.
@@
Random (English-centric) 1,367 587
@@
Random (Intent Classification) 1,820 781
Random (Slot Filling) 1,820 781
Unseen 1,853 101
Cross-Dataset 1,954 224
Table 27: Regressor’s training and test set size on different experimental settings. The total MT experimental records for English-centric and Many-to-Many Languages datasets are 1,954 and 224, respectively. On the other hand, the total experimental records for intent classification and slot filling are both 2,601.
@@
FLoRes Costa-jussà et al. (2022) Muennighoff/flores200 CC-BY-SA 4.0
@@
MASSIVE FitzGerald et al. (2022) https://huggingface.co/datasets/AmazonScience/massive CC-BY 4.0
@@
Figure 5: Detailed results of XGBoost with ProxyLM Ensemble on the M2M100 model under the LOLO setting using the English-centric dataset on MT task from Table 13 per languages.
Figure 6: Detailed results of XGBoost with ProxyLM Ensemble on the LLaMA3 model under the LOLO setting on intent classification from Table 17 per languages.
Figure 7: Detailed results of LGBM with ProxyLM Ensemble on the Aya model under the LOLO setting on slot classification from Table 18 per languages.
Figure 8: Feature importance analysis of XGBoost with ProxyLM Ensemble on MT task using M2M100 model using the English-centric dataset.
Figure 9: Feature importance analysis of XGBoost with ProxyLM Ensemble on MT task using NLLB model using the English-centric dataset.
Figure 10: Feature importance analysis of XGBoost with ProxyLM Ensemble on the Aya-23 on slot filling task.
Figure 11: Feature importance analysis of XGBoost with ProxyLM Ensemble on the LLaMA3 on slot filling task.
Figure 12: Feature importance analysis of XGBoost with ProxyLM Ensemble on the Aya-23 on intent classification task.
Figure 13: Feature importance analysis of XGBoost with ProxyLM Ensemble on the LLaMA3 on intent classification task.